OpenStudy (anonymous):

1+1=??? Creative answers only

OpenStudy (anonymous):

10

OpenStudy (pounce0129):

11

OpenStudy (anonymous):

\[1+1\ge1\]

OpenStudy (anonymous):

\[\frac{ \infty }{ \infty }+\frac{ \infty }{ \infty }\]

OpenStudy (anonymous):

@cosmonautics it could also be -infinity/-infinity lol

OpenStudy (anonymous):

indeed

OpenStudy (3psilon):

1+1 = 10(0) +2(1)

OpenStudy (anonymous):

Yawn....

OpenStudy (cruffo):

0 (mod 2)

OpenStudy (anonymous):

good good answers lol be my fan